# Simulating Relational Event Histories - Why and How

### v2

@@ 3.1 General methodology
    Estimate the parameters $\hat{\beta}_{M_{c}}$ of the candidate model $M_{c}$.
@@
    Generate $L$ sequences $E^{sim}=\{\mathcal{E}_{1},\mathcal{E}_{2}\dots\mathcal{E}_{L}\}$ from the estimates $\hat{\beta}_{M_{c}}$ using the simulation frameworks described in Section 2.
@@
3.2 Example: Assessing goodness-of-fit of an email dataset
We use a subset of publicly available email data from the Enron Corpus (Klimt and Yang,, 2004) to demonstrate the evaluation of goodness-of-fit through simulations. Our purpose is not to provide a comprehensive analysis of goodness-of-fit as this depends on the research question, but to showcase the usefulness and flexibility of simulation techniques for assessing goodness-of-fit for relational event models. The Enron Corpus comprises time-stamped email exchanges between employees of the Enron corporation, a former energy, commodities, and services company. Email datasets are widely used in social network research, including relational event modeling, organization studies, and network science (Mulder and Leenders,, 2019; Quintane et al.,, 2013; Kleinbaum,, 2012; Hardin et al.,, 2015; Estévez-Mujica and Quintane,, 2018). The Enron dataset has been used previously for analysis with relational event models (Perry and Wolfe,, 2013; DuBois et al.,, 2013). For clarity of exposition, we consider a subset of the actors only belonging to either the Legal or Trading department. This subset consists of 60 employees and contains information on the gender, title, seniority, department, and division of these employees. We split the emails with more than one recipient as individual emails between the sender and each of the recipients. Further, we discard emails that contain more than four recipients in order to avoid including broadcast emails that are sent to a large number of receivers as they are covered by very different dynamics than emails exchanged in a smaller group size (Perry and Wolfe,, 2013). A portion of the dataset from January 1, 2001 - August 30, 2001 containing approximately 5000 events is used to illustrate the application of simulation-based goodness-of-fit assessment. For this example, we do not aim to necessarily provide a well-fitting model of the data, but rather will provide a sensible model that might well coincide with an initial model of a researcher who wants to fit a relational event model to this data. This allows us to focus our attention on how simulation could guide the goodness-of-fit process and how it could point the researcher to the strengths and weaknesses of the fitted model.
We compare the goodness-of-fit for two REM models. Both models attempt to capture the communication dynamics in the data through endogenous network effects such as inertia, reciprocity, degree and participating shift effects. See Butts, (2008) for the specification of these statistics. Additionally, the models include exogenous effects such as homophily, i.e. the tendency for actors to interact with other actors sharing the same attributes. The difference between the two models is that one of them accounts for the differences between the impact of events that occurred far in the past versus more recent events. In email interactions, in particular at a corporate organization, the recency of emails (received and sent) plays a crucial role in determining the dynamics of events. We incorporate this into the model following Perry and Wolfe, (2013) by specifying an interval $(\delta_{1},\delta_{2})$ in the past $(t-\delta_{2},t-\delta_{1})$ with $0\leq\delta_{1}<\delta_{2}\leq t$ such that only events that occurred during the interval are counted towards the computation of the statistics. We consider intervals $(0,1),(1,7)$ and $(7,\inf)$ days. Therefore emails in the past day, week, and beyond are considered separately in computing endogenous statistics. Consider $A(t_{1},t_{2})$ to be an $n\times n$ adjacency matrix, where $n$ is the number of actors and each element of the adjacency matrix $a_{ij}(t_{1},t_{2})$ represents the count of events observed for the dyad $(i,j)$ in the time interval $[t_{1},t_{2})$. The definitions of the endogenous statistics and their ‘interval’ counterparts can be formulated as:
inertia: $X(i,j,t)=a_{ij}(0,t)$
interval inertia: $X(i,j,t,\delta_{1},\delta_{2}\mid 0\leq\delta_{2}<\delta_{1}\leq t)=a_{ij}(t-\delta_{1},t-\delta_{2})$
reciprocity: $X(i,j,t)=a_{ji}(0,t)$
interval reciprocity: $X(i,j,t,\delta_{1},\delta_{2}\mid 0\leq\delta_{2}<\delta_{1}\leq t)=a_{ji}(t-\delta_{1},t-\delta_{2})$
out-degree sender: $X(i,j,t)=\sum\limits_{k\in\mathcal{A}/\ i}a_{ij}(0,t)$
interval out-degree sender: $X(i,j,t,\delta_{1},\delta_{2}\mid 0\leq\delta_{2}<\delta_{1}\leq t)=\sum\limits_{k\in\mathcal{A}/\ i}a_{ij}(t-\delta_{1},t-\delta_{2})$
in-degree receiver: $X(i,j,t)=\sum\limits_{k\in\mathcal{A}/\ j}a_{kj}(0,t)$
interval in-degree receiver: $X(i,j,t,\delta_{1},\delta_{2}\mid 0\leq\delta_{2}<\delta_{1}\leq t)=\sum\limits_{k\in\mathcal{A}/\ j}a_{kj}(t-\delta_{1},t-\delta_{2})$
To assess the goodness-of-fit of the models, we first start with the classical measures of goodness-of-fit. The Akaike information criterion (AIC) of the model without interval statistics is 29810 and that of with interval statistics is 28423. Further, the Bayesian information criterion (BIC) of the model with the interval statistics is 28573 and 29921 without. The results suggest that the model with the interval statistics fits the data relatively better than the model without. However, we do not learn about the overall fit and the adequacy of the two models in capturing key characteristics of interest. We also do not gain any insight into how we could potentially improve the fit of the models. A simulation-based goodness-of-fit approach can be utilized in addition to the AIC, BIC measures. To do this we need to select indices that we want the model to reproduce from the observed network. The indices that one would use in a full-fledged empirical study would depend on the research question at hand, the theory being employed, and the baseline against which the researcher wants to compare the model. The indices we have chosen to calculate for this example are meant to provide examples of different kinds of indices one might want to consider in an actual empirical study of this data.
In Figure 1 (1(a)) and (1(b)) we see the in-degree and out-degree distributions of observed and simulated sequences under the two models with and without interval statistics. Both models reproduce the degree distributions in the observed dataset poorly indicating that there might be a mis-fit in the models; therefore, if replicating the degree is important to the researcher, more statistics or attributes need to be included in the model that will attempt to better capture the degree distributions of the data. Figure 1 (1(e)) and (1(f)) report the interevent times for events exchanged in the legal and trading department of the organization. The model with the interval statistics is able to replicate the departmental inter-event times better than the model without interval statistics indicating that including the interval statistics did indeed improve this temporal aspect of the model. Figure 1 (1(c)) shows the proportion of inter-departmental events exchanged in the network per week. We see the model with interval statistics again fits the observed data slightly better than the model without but the real dynamics is is not captured by either model. In Figure 1 (1(d)) we report the proportion of reciprocated ties in the network computed every two weeks. The proportion of reciprocated ties is computed as the proportion of dyads that are symmetric. In our data this includes the dyads with actors who have sent and reciprocated at least one email to each other in the past two weeks. Both models replicate the observed reciprocated ties poorly as they generate inter-departmental events that are fairly stable. This is expected as neither the parameters nor the attributes change over time in the fitted model which generates a stable sequence of interactions over time. Any researcher with data that changes considerably over time, should bear in mind to include time varying effects or attributes or other temporal aspects that may account for the deviation from the equilibrium. Thus we conclude that simulation-based goodness-of-fit assessment provides a fine-grained understanding of the fit of relational event models and allows researchers to evaluate potential areas where the model specification could be improved further.
(a) In-degree distribution
(b) Out-degree distribution
(c) Proportion of Inter-department Events
(d) Proportion of Reciprocated Ties
(e) Interevent Time - Legal Department
(f) Interevent Time - Trading Department
@@ 3.2 Example: Assessing goodness-of-fit of an email dataset
Figure 1: Goodness-of-fit comparison. Each figure plots the values of the fit indices over 50 simulated sequences for each model.
@@ 4 Developing theories using simulations
The second important application of the relational event simulations we present here involves the development of social theories using simulations. Many social phenomenon arise from repeated interactions among individuals in a social network over time. Undoubtedly, researchers continue to gain insights into the drivers of repeated social interactions through empirical studies. The relational event model itself has contributed greatly towards modelling the data arising out of such empirical research. While this is useful for estimating the magnitude of network effects or confirming differences between different groups in specific empirical settings, fitted REMs are typically based on theoretical ideas of the researcher and contribute to testing existing theories rather than being employed as a way to develop or refine (new) theory. Simulations can provide a powerful and flexible tool for (further) developing or expanding theories (Davis et al.,, 2007) regarding why, how, and when interactions occur in a social network. The proposed relational event simulation frameworks provide flexibility to combine multiple network effects such as transitivity, homophily, or preferential attachment in one theoretical model, allowing researchers to test and develop richer theories about social interaction dynamics. Below we first describe how we can (further) develop social network theories using relational event simulations, and subsequently, we present an extensive application of the methodology to better understand group formation using optimal distinctive theory.
@@ 4.2 Group formation using Optimal Distinctiveness Theory
In our simulations, we assume that all actors have the same value of optimal distinctiveness $d^{*}$. We simulate networks consisting of 30 actors with a binary attribute that represents their fixed identity $z=\{1,0\}$ for varying values of $d^{*}$. The proportion of actors with each attribute is equal. To identify emergent groups formed as a result of the simulations, we use the Louvain community detection algorithm Blondel et al., (2008). This method of community detection essentially partitions nodes based on a greedy modularity optimisation. The groups identified from this method are considered emergent because they arise from interactions and decisions made by actors on the actor level which lead to emergent group formation on a network level. Figure 2 depicts the network plots for simulated sequences with different $d^{*}$ values. When $d^{*}=0$ as illustrated in Figure 2 (2(a)), actors in the simulation strongly prefer homophilic interactions, resulting in groups containing nodes of one attribute type. Actors with opposing attributes tend to cluster together rather than interact with actors of a different attribute.
@@ 5.1 Types of network interventions
5.2 Evaluating the Persistence of Intervention Outcomes
In this example, our aim is to explore the longevity of outcomes of different interventions through simulations. With simulations, it is possible to assess how a network would react to competing interventions and assists in improving intended network interventions towards a desirable outcome. The intervention in this example is intended to increase the inter-department collaboration between employees of two departments in an organization. The number of inter-departmental emails is used as a proxy for the collaboration across departments. This intervention example is a simplified version of an actual intervention study that we conducted for the R&D department of a large consumer goods company in Western Europe. To increase the inter-department communication, the organisation intended to intervene by providing the employees with incentives to increase their inter-department communication (e.g., by organising new multi-department projects and organising joint meetings). This intervention is modelled in our example by temporarily increasing the departmental heterophily effect $\beta_{het}$ in the model. Our interest is in evaluating what the effect is of the intervention after it ends and how long the increased heterophily effect remains in the network after the intervention has ended and the network has returned to its “normal” dynamic. Our assumption is that the intervention by itself will not structurally affect employee interaction norms and preferences. Therefore, after the intervention, we assume that the “pre-intervention” model is again applicable to the interaction among the employees (but the interaction histories will have been altered by the interventions).
Before testing the interventions using simulations, we first estimate the coefficients of a model based on interactions before the intervention, for a period spanning 243 days using a standard REM. The REM model contains various endogenous effects (represented by $\beta_{endo}$) such as inertia, reciprocity, participation shifts and degree effects. Exogenous effects (represented by $\beta_{exo}$) for the gender, seniority and sub-domain are also included. An additional effect for inter-departmental heterophily ($\beta_{het}$ is also included and is the locus for the intervention.
The model fits the data reasonably well and we assume that it captures the way in which the company’s employees tend to interact well. In the following, we will compare nine versions of the intervention to increase the inter-departmental homophily (by increasing $\beta_{het}$). The interventions differ in the length of time they run and in the strength of the intervention and we aim to evaluate the persistence of the intervention effects after the intervention has been carried out. Figure 4 graphically shows two examples. Intervention I lasts from $t_{1}$ to $t_{2}$ and for this period we increase $\beta_{het}$ (the parameter for the heterophily effect) to $\beta_{2}$. Alternatively, intervention II runs longer than intervention I (from $t_{1}$ to $t_{2}^{\prime}$) and has a different heterophily parameter $\beta_{het}=\beta_{3}$.
@@ 5.2 Evaluating the Persistence of Intervention Outcomes
Using the pre-intervention sequence as the initialization, we simulate the relational events in the network for the length of time the intervention runs and following the intervention, we further simulate the relational sequence for 50 weeks after each intervention. We consider three durations of intervention $\{28,42,56\}$ days and three inter-departmental heterophily effect sizes $\{0,0.25,0.5\}$. Therefore, we generate relational event sequences under $3\times 3=9$ different interventions. Since the simulations have a stochastic component, we generate 50 relational event sequences for each intervention and compute the proportion of interdepartmental events in each. For each intervention, we report the median proportion in the 50 simulated sequences in Figure 5. The proportion of inter-department events increased most for interventions held for the longest period and for the strongest effect size. The effects of the interventions on the inter-departmental communication lasted 38 days for the case of $\beta_{het}=0$, 91 days for the case of $\beta_{het}=0.25$ and 208 days when $\beta_{het}=0.5$ when the duration of the intervention period was 28 days. Similarly, the effects of the intervention last 61 days for $\beta_{het}=0$, 247 days for $\beta_{het}=0.25$, and 283 days for $\beta_{het}=0.5$ when intervention duration was 42 days. Further, the impact of the intervention lasted 72 days for $\beta_{het}=0$, 307 days for $\beta_{het}=0.25$, and persist beyond our simulation period when $\beta_{het}=0.5$ when intervention duration was 56 days. We conclude that relational event simulations can provide a useful tool in planning and designing network interventions as well as to evaluate the persistence of interventions outcomes over time.

### v4

@@ 3.1 General methodology
    Estimate the parameters $\hat{\boldsymbol{\beta}}_{M_{c}}$ of the candidate model $M_{c}$.
@@
    Generate $L$ sequences $E^{sim}=\{\mathcal{E}_{1},\mathcal{E}_{2}\dots\mathcal{E}_{L}\}$ from the estimates $\hat{\boldsymbol{\beta}}_{M_{c}}$ using the simulation frameworks described in Section 2.
@@
3.2 Example: Assessing goodness-of-fit
To demonstrate the evaluation of goodness-of-fit through simulations, we evaluate the fit of a model presented in a recent study by Gravel et al., (2023). The study focuses on the dynamics of inter-gang violence in Los Angeles using a relational event modeling approach. The relational event history comprises of violent incidents involving 33 gangs spanning over three years. The authors examine various mechanisms that drive the continuation of violence in this network. The authors focus on how gang characteristics, spatial proximity, enduring rivalries, and how the dynamics of past conflicts influence future violent events. The authors argue that while retaliation is a significant mechanism, other factors like established rivalries and inertia play crucial roles in driving future violence. The authors of the article attempt to assess the goodness of fit of their model by checking how well the fitted model can predict the observed events in-sample: For every observed event, it was checked how likely it was to be observed given the (endogenous) predictor variables at that point. Even though this analysis indicated that the fitted model has a high predictive accuracy, this assessment is unable to answer whether the entire sequence of relational events is plausible under the fitted model when it would be used to generate entire event sequences “from scratch”. That is the purpose of the current simulation analysis. Moreover, following step 1 of the goodness of fit assessment procedure presented in Section 3.1, we shall investigate the fit of various data characteristics on gang violence including gang size, inter-dyadic distance, geodesic hop-pah lengths, degree statistics, and inter-event times between rival gangs.
(a) Inter-gang size event density
(b) Inter-dyadic distance event density
(c) Temporal geodesic hop-path length distribution
(d) Participating shift counts
(e) Out-degree distribution
(f) In-degree distribution
(g) Interevent time - rivals
(h) Interevent reciprocation time - rivals
@@ 3.2 Example: Assessing goodness-of-fit
Figure 1: Goodness-of-fit comparison. Each figure plots the values of the fit indices over 100 simulated sequences and the observed data. In figures (a-f) the red crosses represent the observed data and the boxplots depict the distribution of the fit indices across simulated networks. In figures (g-h) the dotted red line represents the observed data.
To assess the goodness of fit for the relational event models presented in this study, we simulate from the fitted relational event models (see Appendix B for R-code) using the maximum likelihood estimates presented in Table 2 of the study. The plots in Figure 1 illustrate the goodness-of-fit for the model, comparing the characteristics of simulated networks against actual observed data in the context of street gang violence. Each subplot highlights a different network characteristic derived from violent conflicts between gangs. Figure 1(1(a)) shows the density of events categorized by size of participating gangs in the dyads. In the context of the paper, understanding these interactions is crucial because larger gangs might have different resources, territorial ambitions, and levels of influence, which can affect their interaction patterns with smaller gangs. The observed data (red crosses) are relatively well-aligned with the median of the simulated distributions, implying that the model captures the frequency of events across gang-size effectively. There is a slight misfit involving dyads with a larger gang size which might necessitate a need for an additional variable to effectively reproduce the interaction patterns of larger sized gangs. In Figure 1(1(b)) the density of events is plotted against the geographic distance between dyads. The article highlights the importance of spatial proximity in the dynamics of gang violence, making it vital for the model to accurately simulate these interactions. The model appears to accurately capture the decline in event density as distance increases. The model slightly underestimates the frequency of localized conflicts which are more frequent and significant within closer geographical proximity. The Figure 1(1(c)) examines the distribution of temporal path lengths with shorter paths indicating more direct interactions over time between gangs (with lesser temporally respecting hops). The model reproduces the geodesic hop-path length distribution across the dyads in the network reasonably well. This fit index helps understand how violence spreads over time through the network, testing theories of violence contagion and diffusion within gang dynamics.
Participating shift counts of dyadic turn-taking events (Gibson,, 2003) are depicted in 1(1(d)). While, AB-BA and AB-BY type participating shifts were included in the model, the results suggest that the model is also able to reproduce the AB-AY, AB-XA and AB-XB turn taking events well. The out-degree distribution represents the number of conflicts initiated by each gang is depicted in 1(1(e)). The observed data mostly aligns with the simulated medians for lower out-degrees but tends to fall outside the interquartile range for higher degrees. This could be due to the model not accounting for potential external factors that might influence a gang’s propensity for aggressive conflict initiation. In addition we also look at the in-degree distribution of the data in Figure 1(1(f)) which suggest an adequate fit.
The last two figures 1(1(g),1(h)) show the distribution of interevent events categorized by inertia and reciprocation among rival gangs. These timing-based indices test the model’s ability to quantify the rapidity and persistence of violence between rival gangs. It is vital for investigating how well the model replicates violence escalation and retaliation cycles, which are key aspects of gang dynamics described in the paper In the figure, the observed data, is represented by the red dotted line. Both plots demonstrate that conflicts between rival gangs are characterized by rapid sequences of events, with a significant number of actions occurring within a very short time frame following an initial event. Although, the plot for reciprocation times between rival gangs has a somewhat less steep decline compared to the interevent times for rivals. This suggests that while immediate responses are common, there’s slightly more variation in the timing of reciprocations, possibly due to differing strategic considerations or external factors influencing the timing of a gang’s response. The simulated results indicate an underestimation at very short inter-event times, suggesting the model might not fully capture the quickest escalations in gang conflict. Refining the model by including an interaction term between the recent reciprocation and recent target persistence and rivalry might address the misfit. Additionally, short term-memory effects (Brandes et al.,, 2009) that allow for more rapid changes in event probabilities immediately following a conflict event might allow the model to address the misfit.
Evaluating the goodness-of-fit using simulations has provided us with a more nuanced understanding of how well the model fits and highlighted potential areas of misfit compared to predictive performance or likelihood based measures. Overall, while the model provides a good general fit to the observed data across several key indices, the detailed analysis provided insights into where the model falls short in capturing the dynamics of gang conflicts, such as the inter-event times between rival gangs. Understanding and refining these aspects of the model could lead to better predictions of gang conflict dynamics and more effective strategies for law enforcement and community interventions aimed at reducing the frequency and severity of these conflicts.
@@ 4 Developing theories using simulations
The second important application of the relational event simulation techniques involves the development of social theories. Many social phenomenon arise from repeated interactions among individuals in a social network over time. Undoubtedly, researchers continue to gain insights into the drivers of repeated social interactions through empirical studies. The relational event model itself has contributed greatly towards modelling the data arising out of such empirical research. While this is useful for estimating the magnitude of network effects or confirming differences between different groups in specific empirical settings, fitted REMs are typically based on theoretical ideas of the researcher and contribute to testing existing theories rather than being employed as a way to develop or refine (new) theory. Simulations can provide a powerful and flexible tool for (further) developing or expanding theories (Davis et al.,, 2007) regarding why, how, and when interactions occur in a social network. The proposed relational event simulation frameworks provide flexibility to combine multiple network effects such as transitivity, homophily, or preferential attachment in one theoretical model, allowing researchers to test and develop richer theories about social interaction dynamics. Below we first describe how we can (further) develop social network theories using relational event simulations, and subsequently, we present an extensive application of the methodology to better understand group formation using optimal distinctive theory.
@@ 4.2 Group formation using Optimal Distinctiveness Theory
In our simulations, we assume that all actors have the same value of optimal distinctiveness $d^{*}$. We simulate networks consisting of 30 actors with a binary attribute that represents their fixed identity $z=\{1,0\}$ for varying values of $d^{*}$. The proportion of actors with each attribute is equal. To identify emergent groups formed as a result of the simulations, we use the Louvain community detection algorithm (Blondel et al.,, 2008). This method of community detection essentially partitions nodes based on a greedy modularity optimisation. The groups identified from this method are considered emergent because they arise from interactions and decisions made by actors on the actor level which lead to emergent group formation on a network level. Figure 2 depicts the network plots for simulated sequences with different $d^{*}$ values. When $d^{*}=0$ as illustrated in Figure 2 (2(a)), actors in the simulation strongly prefer homophilic interactions, resulting in groups containing nodes of one attribute type. Actors with opposing attributes tend to cluster together rather than interact with actors of a different attribute.
@@ 5.1 Types of network interventions
5.2 Evaluating the Persistence of Intervention Outcomes
In this example, our aim is to explore the longevity of outcomes of different interventions through simulations. With simulations, it is possible to assess how a network would react to competing interventions and assists in improving intended network interventions towards a desirable outcome. The intervention in this example is intended to increase the inter-department collaboration between employees of two departments in an organization. The number of inter-departmental emails is used as a proxy for the collaboration across departments. This intervention example is a simplified version of an actual intervention study that we conducted for the R&D department of a large consumer goods company in Western Europe. To increase the inter-department communication, the organisation intended to intervene by providing the employees with incentives to increase their inter-department communication (e.g., by organising new multi-department projects and organising joint meetings). This intervention is modelled in our example by temporarily increasing the departmental heterophily effect $\boldsymbol{\beta}(t)_{het}$ in the model. Our interest is in evaluating what the effect is of the intervention after it ends and how long the increased heterophily effect remains in the network after the intervention has ended and the network has returned to its “normal” dynamic. Our assumption is that the intervention by itself will not structurally affect employee interaction norms and preferences. Therefore, after the intervention, we assume that the “pre-intervention” model is again applicable to the interaction among the employees (but the interaction histories will have been altered by the interventions).
Before testing the interventions using simulations, we first estimate the coefficients of a model based on interactions before the intervention, for a period spanning 243 days using a standard REM. The REM model contains various endogenous effects (represented by $\boldsymbol{\beta}(t)_{endo}$) such as inertia, reciprocity, participation shifts and degree effects. Exogenous effects (represented by $\boldsymbol{\beta}(t)_{exo}$) for the gender, seniority and sub-domain are also included. An additional effect for inter-departmental heterophily ($\boldsymbol{\beta}(t)_{het}$ is also included and is the locus for the intervention.
The model fits the data reasonably well and we assume that it captures the way in which the company’s employees tend to interact well. In the following, we will compare nine versions of the intervention to increase the inter-departmental homophily (by increasing $\boldsymbol{\beta}(t)_{het}$). The interventions differ in the length of time they run and in the strength of the intervention and we aim to evaluate the persistence of the intervention effects after the intervention has been carried out. Figure 4 graphically shows two examples. Intervention I lasts from $t_{1}$ to $t_{2}$ and for this period we increase $\boldsymbol{\beta}(t)_{het}$ (the parameter for the heterophily effect) to $\boldsymbol{\beta}(t)_{2}$. Alternatively, intervention II runs longer than intervention I (from $t_{1}$ to $t_{2}^{\prime}$) and has a different heterophily parameter $\boldsymbol{\beta}(t)_{het}=\boldsymbol{\beta}(t)_{3}$.
@@ 5.2 Evaluating the Persistence of Intervention Outcomes
Using the pre-intervention sequence as the initialization, we simulate the relational events in the network for the length of time the intervention runs and following the intervention, we further simulate the relational sequence for 50 weeks after each intervention. We consider three durations of intervention $\{28,42,56\}$ days and three inter-departmental heterophily effect sizes $\{0,0.25,0.5\}$. Therefore, we generate relational event sequences under $3\times 3=9$ different interventions. Since the simulations have a stochastic component, we generate 50 relational event sequences for each intervention and compute the proportion of interdepartmental events in each. For each intervention, we report the median proportion in the 50 simulated sequences in Figure 5. The proportion of inter-department events increased most for interventions held for the longest period and for the strongest effect size. The effects of the interventions on the inter-departmental communication lasted 38 days for the case of $\boldsymbol{\beta}(t)_{het}=0$, 91 days for the case of $\boldsymbol{\beta}(t)_{het}=0.25$ and 208 days when $\boldsymbol{\beta}(t)_{het}=0.5$ when the duration of the intervention period was 28 days. Similarly, the effects of the intervention last 61 days for $\boldsymbol{\beta}(t)_{het}=0$, 247 days for $\boldsymbol{\beta}(t)_{het}=0.25$, and 283 days for $\boldsymbol{\beta}(t)_{het}=0.5$ when intervention duration was 42 days. Further, the impact of the intervention lasted 72 days for $\boldsymbol{\beta}(t)_{het}=0$, 307 days for $\boldsymbol{\beta}(t)_{het}=0.25$, and persist beyond our simulation period when $\boldsymbol{\beta}(t)_{het}=0.5$ when intervention duration was 56 days. We conclude that relational event simulations can provide a useful tool in planning and designing network interventions as well as to evaluate the persistence of interventions outcomes over time.